# Learnability is a Compact Property

### v1

@@ 3.2 Improper Loss Functions
Let $\mathcal{Y}$ be a metric space and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Fix $\xi:\mathbb{N}\to\mathbb{R}_{\geq 0}$, and suppose that for any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\xi_{\mathcal{H}^{\prime}}\leq\xi$. Then $\xi_{\mathcal{H}}\leq 2\cdot\xi$.
@@
The proof of Theorem 3.8 is deferred to Appendix B.2, but let us briefly sketch the main idea. Fix $n\in\mathbb{N}$, $S\in\mathcal{X}^{n}$, and set $\epsilon=\xi(n)$. Consider again the collection of functions $R=\mathcal{H}|_{S}$ and variables $L=\bigcup_{S^{\prime}\subseteq S,|S^{\prime}|=n-1}\mathcal{H}|_{S^{\prime}}$, as described in the proof of Theorem 3.6. By the premise of the theorem, for any finite subset $R^{\prime}\subseteq R$, there exists an assignment of variables $L\to\mathcal{Y}$ which maintains all functions in $R^{\prime}$ below $\epsilon$. Each such assignment induces an apportionment of error to each function in $R^{\prime}$, i.e., a vector of length $n$ with positive entries summing to $\epsilon$. For $r\in R^{\prime}$ depending upon variables $\ell_{1},\ldots,\ell_{n}$, this apportionment tracks the contribution of each $\ell_{i}$ to the output of $r$. The central technical step of the proof is to demonstrate that one can assign apportionments to each node $r\in R$ such that any finite subset of the apportionments can be satisfied by an assignment of variables $L\to\mathcal{Y}$. Then let $\ell=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$ be a variable. We assign $\ell$ to the value $\hat{y}$ such that the function $(y_{1},\ldots,y_{i},\hat{y},y_{i+1},\ldots,y_{n})$ has minimal budget apportioned to $\ell$, among all such $\hat{y}$. From an invokation of the triangle inequality, this learner at most doubles the output of any $r\in R$.
@@
Recall from Section 3.1 that proper metric spaces are sufficiently expressive to describe many of the most frequently studied label spaces, including $\mathbb{R}^{n}$ (equipped with any norm) and its closed subsets. What, then, is a typical example of a label space which fails to be improper? Perhaps the most natural example is multiclass classification over infinite label sets, i.e., $\mathcal{Y}$ equipped with the discrete metric $\ell_{0-1}(y,y^{\prime})=[y\neq y^{\prime}]$. We will now demonstrate, however, that the particular structure of multiclass classification can be exploited to recover an exact compactness result in the style of Theorem 3.6. Notably, we do so by invoking M. Hall’s classic matching theorem for infinite graphs, which for good measure we show to be a special case of our Theorem 3.3.
@@
M. Hall’s theorem states that an infinite bipartite graph $G$ is $R$-matchable if and only if it is finitely $R$-matchable, provided that all nodes in $R$ have finite degree. Before proving M. Hall’s theorem by way of Theorem 3.3, it will be useful to establish an intermediate lemma: informally, all nodes in $G$ can be taken to have finite degree, not merely those in $R$.
@@
The proof of Lemma 3.11 is deferred to Appendix B.3, but its intuition is fairly simple: by P. Hall’s theorem, $G$ is finitely $R$-matchable precisely when Hall’s condition holds, i.e., $|N(R^{\prime})|\geq|R^{\prime}|$ for all finite $R^{\prime}\subseteq R$ [23]. Thus any $\ell\in L$ which is not incident to a Hall blocking set can be removed from $G$ while preserving Hall’s condition and finite $R$-matchability. Proceeding in this way, nodes can be removed until each remaining $\ell\in L$ is contained in a Hall blocking set $R^{\prime}_{\ell}$. At this point, $\ell$’s incident edges can be safely restricted to those which are incident with $R^{\prime}_{\ell}$, a finite set.
@@
The forward direction is clear. For the reverse, suppose $G$ is finitely $R$-matchable. Then we may assume as a consequence of Lemma 3.11 that the nodes in $L$ have finite degree as well. Let us think of each node $\ell\in L$ as a variable residing in the discrete metric space on its neigbhors. We will also think of each node $r\in R$ as a function of its neighbors, which outputs the number of neighbors that have not been assigned to $r$ itself. Note that the discrete metric space on finitely many elements is proper, and furthermore that any function out of such a space is automatically proper. Then invoke Theorem 3.3 with $\epsilon(r)=1-\frac{1}{\deg(r)}$ to complete the proof. (See Remark 3.5.) ∎
@@
Let $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a classification problem. Then the following are equivalent:
@@
    $\mathcal{H}$ is learnable in the realizable case with transductive sample function $m:(0,1)\to\mathbb{N}$.
@@
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable in the realizable case with transductive sample function $m:(0,1)\to\mathbb{N}$.
@@ 3.3 Agnostic Learning
Our discussion thus far has restricted attention to learning in the realizable case: what can be said of the agnostic case? In short, all results from Sections 3.1 and 3.2 can be claimed for the agnostic case (with nearly identical proofs), with the exception of Theorem 3.8. To begin, let us briefly review transductive learning in the agnostic case. See [7] for a more detailed discussion.
@@
    One labeled datapoint $(x_{i},y_{i})$ is selected uniformly at random from $S$. The remaining datapoints $S_{-i}$ are displayed to the learner.
@@
Notably, tranductive learning in the agnostic case differs from the realizable case in that the adversary is no longer restricted to label the datapoints in $S$ using a hypothesis $\mathcal{H}$. To compensate for the increased difficulty, and in accordance with the PAC definition of agnostic learning, a learner is only judged relative to best-in-class performance across $\mathcal{H}$. Formally, the loss of a learner on an instance $S\in(\mathcal{X}\times\mathcal{Y})^{n}$ is defined
@@
Furthermore, one can use nearly identical reasoning as in the proof of Theorem 3.6 to see that agnostic transductive learning is described by a similar system of variables $L$ and functions $R$. In particular, set $R=\mathcal{Y}^{n}$ and let $L\subseteq(\mathcal{Y}\cup\{?\})^{n}$ contain all sequences with exactly one ``?"``?"``?"` ` ? ". Then a function $r=(y_{1},\ldots,y_{n})$ depends upon the variables $\{\ell_{i}\}_{i\in[n]}$, where $\ell_{i}=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$ and
@@
Now, as in the realizable case, a learner $A$ corresponds precisely to an assignment of each variable $\ell\in L$ to a value in $\mathcal{Y}$, and $A$ incurs agnostic transductive error at most $\epsilon$ if and only if the outputs of all nodes in $R$ are maintained below $\epsilon$. When $\mathcal{Y}$ is a proper metric space, exact compactness of sample complexity thus comes as an immediate consequence of Theorem 3.3 and our preceding discussion. Furthermore, when $\mathcal{Y}$ bears a discrete metric, learning reduces to an assignment problem in graphs, and exact compactness follows from a straightforward splitting argument applied to M. Hall’s matching theorem (as in Corollary 3.13). We thus have the following theorem.
@@
Let $\mathcal{Y}$ a label set with the structure of a proper metric space or a discrete metric space, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Then the following conditions are equivalent:
@@
    $\mathcal{H}$ is learnable in the agnostic case with transductive sample function $m:(0,1)\to\mathbb{N}$.
@@
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable in the agnostic case with transductive sample function $m:(0,1)\to\mathbb{N}$.
@@
Regarding improper metric spaces, note that our lower bound from Theorem 3.7 transfers directly to the agnostic case, as it established for a hypothesis class for which agnostic learning is precisely as difficult as realizable learning. We conjecture that larger differences in such error rates — perhaps of arbitrarily large ratio — are possible for the agnostic case.
Let us mention briefly that the connection between transductive learning and PAC learning may not be as tight in the agnostic case as in the realizable case. To our knowledge, there is no result bridging transductive and PAC sample complexities in the style of Lemma 2.6 for the agnostic case.

### v2

@@ 3.2 Realizable Learning With Improper Loss Functions
There exists a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, metric loss function $d$ on $\mathcal{Y}$, and $n\in\mathbb{N}$ such that for any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\xi_{\mathcal{H}}(n)\geq 2\cdot\xi_{\mathcal{H}^{\prime}}(n)$.
Let us describe the main idea of Theorem 3.8, whose proof is deferred to Appendix B.1. The crucial step lies in the creation of the label space $\mathcal{Y}=R\cup S$, where $R$ is an infinite set whose points are all distance 2 apart, and $S$ is an infinite set whose elements are indexed by the finite subsets of $R$, e.g., as in $s_{R^{\prime}}\in S$ for finite $R^{\prime}\subseteq R$. For all such $R^{\prime}$, define $s_{R^{\prime}}$ to be distance 1 from the elements of $R^{\prime}$, distance 2 from the other elements of $R$, and distance 1 from all other points in $S$. Then $\mathcal{Y}$ indeed forms a metric space, and it is straightforward to see that, for instance, the class of all functions from a one-element set to $\mathcal{Y}$ is more difficult to learn than its finite projections (equivalently, finite subsets).
We now prove a matching upper bound to Theorem 3.8, demonstrating that a factor of 2 is the greatest possible gap between the error rate of $\mathcal{H}$ and its projections when the loss function is a metric.
Theorem 3.9.
Let $\mathcal{Y}$ be a label set with a metric loss function and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Fix $\xi:\mathbb{N}\to\mathbb{R}_{\geq 0}$, and suppose that for any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ has transductive error rate $\xi_{\mathcal{H}^{\prime}}\leq\xi$. Then $\mathcal{H}$ has transductive error rate $\xi_{\mathcal{H}}\leq 2\cdot\xi$.
@@
The proof of Theorem 3.9 is deferred to Appendix B.2, but let us briefly sketch the main idea. Fix $n\in\mathbb{N}$, $S\in\mathcal{X}^{n}$, and set $\epsilon=\xi(n)$. Consider again the collection of functions $R=\mathcal{H}|_{S}$ and variables $L=\bigcup_{S^{\prime}\subseteq S,|S^{\prime}|=n-1}\mathcal{H}|_{S^{\prime}}$, as described in the proof of Theorem 3.6. By the premise of the theorem, for any finite subset $R^{\prime}\subseteq R$, there exists an assignment of variables $L\to\mathcal{Y}$ which maintains all functions in $R^{\prime}$ below $\epsilon$. Each such assignment induces an apportionment of error to each function in $R^{\prime}$, i.e., a vector of length $n$ with positive entries summing to $\epsilon$. For $r\in R^{\prime}$ depending upon variables $\ell_{1},\ldots,\ell_{n}$, this apportionment tracks the contribution of each $\ell_{i}$ to the output of $r$. The central technical step of the proof is to demonstrate that one can assign apportionments to each node $r\in R$ such that any finite subset of the apportionments can be satisfied by an assignment of variables $L\to\mathcal{Y}$. Then let $\ell=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$ be a variable. We assign $\ell$ to the value $\hat{y}$ such that the function $(y_{1},\ldots,y_{i},\hat{y},y_{i+1},\ldots,y_{n})$ has minimal budget apportioned to $\ell$, among all such $\hat{y}$. From an invocation of the triangle inequality, this learner at most doubles the output of any $r\in R$.
@@
Recall from Section 3.1 that proper metric spaces are sufficiently expressive to describe many of the most frequently studied label spaces, including $\mathbb{R}^{d}$ (equipped with any norm) and its closed subsets. What, then, is a typical example of a label space which fails to be proper? Perhaps the most natural example is multiclass classification over infinite label sets, i.e., $\mathcal{Y}$ equipped with the discrete metric $\ell_{0-1}(y,y^{\prime})=[y\neq y^{\prime}]$. We will now demonstrate, however, that the particular structure of multiclass classification can be exploited to recover an exact compactness result in the style of Theorem 3.6. Notably, we do so by invoking M. Hall’s classic matching theorem for infinite graphs, which for good measure we show to be a special case of our Theorem 3.3.
@@
M. Hall’s theorem states that an infinite bipartite graph $G$ is $R$-matchable if and only if it is finitely $R$-matchable, provided that all nodes in $R$ have finite degree. Before proving M. Hall’s theorem by way of Theorem 3.3, we establish an intermediate lemma: informally, all nodes in $G$ can be taken to have finite degree, not merely those in $R$.
@@
The proof of Lemma 3.12 is deferred to Appendix B.3, but its intuition is fairly simple: by P. Hall’s theorem, $G$ is finitely $R$-matchable precisely when Hall’s condition holds, i.e., $|N(R^{\prime})|\geq|R^{\prime}|$ for all finite $R^{\prime}\subseteq R$ (Hall, 1935). Thus any $\ell\in L$ which is not incident to a Hall blocking set can be removed from $G$ while preserving Hall’s condition and finite $R$-matchability. Proceeding in this way, nodes can be removed until each remaining $\ell\in L$ is contained in a Hall blocking set $R^{\prime}_{\ell}$. At this point, $\ell$’s incident edges can be safely restricted to those which are incident with $R^{\prime}_{\ell}$, a finite set.
@@
The forward direction is clear. For the reverse, suppose $G$ is finitely $R$-matchable. Then we may assume as a consequence of Lemma 3.12 that the nodes in $L$ have finite degree as well. Let us think of each node $\ell\in L$ as a variable residing in the discrete metric space on its neighbors. We will also think of each node $r\in R$ as a function of its neighbors, which outputs the number of neighbors that have not been assigned to $r$ itself. Note that the discrete metric space on finitely many elements is proper, and furthermore that any function out of such a space is automatically proper. Then invoke Theorem 3.3 with $\epsilon(r)=1-\frac{1}{\deg(r)}$ to complete the proof. (See Remark 3.5.) ∎
@@
Let $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a classification problem, i.e., employing the 0-1 loss function. Then the following are equivalent for any $m\colon\mathbb{R}_{>0}\to\mathbb{N}$:
@@
    $\mathcal{H}$ is learnable in the realizable case with transductive sample function $m$.
@@
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable in the realizable case with transductive sample function $m$.
@@ 3.3 Agnostic Learning
Our discussion thus far has restricted attention to realizable learning: what can be said of the agnostic case? In short, all results from Sections 3.1 and 3.2 can be claimed for agnostic learning (with nearly identical proofs), with the exception of Theorem 3.9. To begin, let us briefly review transductive learning in the agnostic case. See Asilis et al. (2024) or Dughmi et al. (2024) for further detail.
@@
    The unlabeled datapoints in $S$ are all revealed to the learner.
@@
    One labeled datapoint $(x_{i},y_{i})$ is selected uniformly at random from $S$. The remaining labeled datapoints $S_{-i}$ are displayed to the learner.
  • 4.
@@
Notably, transductive learning in the agnostic case differs from the realizable case in that the adversary is no longer restricted to label the datapoints in $S$ using a hypothesis $\mathcal{H}$. To compensate for the increased difficulty, and in accordance with the PAC definition of agnostic learning, a learner is only judged relative to best-in-class performance across $\mathcal{H}$. Formally,
@@
Furthermore, one can use nearly identical reasoning as in the proofs of Theorems 3.6 and 3.7 to see that agnostic transductive learning is described by a system of variables $L$ and functions $R$. In particular, set $R=\mathcal{Y}^{n}$ and let $L\subseteq(\mathcal{Y}\cup\{?\})^{n}$ contain all sequences with exactly one ``?"``?"``?"` ` ? ". Then a function $r=(y_{1},\ldots,y_{n})$ depends upon the variables $\{\ell_{i}\}_{i\in[n]}$, where $\ell_{i}=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$ and
@@
Now, as in the realizable case, a learner $A$ corresponds precisely to an assignment of each variable $\ell\in L$ to a value in $\mathcal{Y}$, and $A$ incurs agnostic transductive error at most $\epsilon$ if and only if the outputs of all nodes in $R$ are maintained below $\epsilon$. Under the conditions of Theorems 3.6 or 3.7, exact compactness of sample complexity thus comes as an immediate consequence of Theorem 3.3 and our preceding discussion. Furthermore, when $\mathcal{Y}$ bears a discrete metric, learning reduces to an assignment problem in graphs, and exact compactness follows from a straightforward splitting argument applied to M. Hall’s matching theorem (as in Corollary 3.14). We thus have the following theorem.
@@
Let $\mathcal{Y}$ and the loss function satisfy the conditions of Theorem 3.6, Theorem 3.7, or Corollary 3.14. Then the following conditions are equivalent for any $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ and $m\colon\mathbb{R}_{>0}\to\mathbb{N}$:
@@
    $\mathcal{H}$ is learnable in the agnostic case with transductive sample function $m$.
@@
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable in the agnostic case with transductive sample function $m$.
@@
Regarding improper metric losses, note that our lower bound from Theorem 3.8 transfers directly to the agnostic case, as it established for a hypothesis class for which agnostic learning is precisely as difficult as realizable learning. We conjecture that larger differences in such error rates — perhaps of arbitrarily large ratio — are possible for the agnostic case.
3.4 PAC Learning
Though our results have thus far been phrased in the language of transductive learning, we now demonstrate that they may be easily extended (in an approximate manner) to Valiant’s celebrated PAC model (Valiant, 1984). The PAC model makes use of probability measures $D$ over $\mathcal{X}\times\mathcal{Y}$, for which the true error incurred by a predictor $h$ is defined as $L_{D}(h)=\operatorname*{\mathbb{E}}_{(x,y)\sim D}\ell(h(x),y)$.
Definition 3.17.
$L_{D}(A(S))\leq\inf_{\mathcal{H}}L_{D}(h)+\epsilon.$
Agnostic PAC learning refers to the case in which $\mathbb{D}$ consists of all measures over $\mathcal{X}\times\mathcal{Y}$, and realizable PAC learning to the case in which $\mathbb{D}=\{D:\min_{\mathcal{H}}L_{D}(h)=0\}$.
Definition 3.18.
The sample complexity of a learner $A$ with respect to a hypothesis class $\mathcal{H}$, $m_{\operatorname{PAC},A}:(0,1)^{2}\to\mathbb{N}$, is the minimal sample function it attains as a learner for $\mathcal{H}$. The sample complexity of a class $\mathcal{H}$ is the pointwise minimal sample complexity attained by any of its learners, i.e., $m_{\operatorname{PAC},\mathcal{H}}(\epsilon,\delta)=\min_{A}m_{\operatorname{PAC},A}(\epsilon,\delta)$.
As previously mentioned, transductive learning bears a close connection to PAC learning: see Asilis et al. (2024) and Dughmi et al. (2024) for further detail on their approximate equivalence.
Lemma 3.19 (Asilis et al. (2024, Proposition 3.6)).
Let $\mathcal{X}$ be a domain, $\mathcal{Y}$ a label set, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Fix a loss function taking values in $[0,1]$. Then the following inequality holds for all $\epsilon,\delta\in(0,1)$ and the constant $e\approx 2.718$:
$m_{\operatorname{Trans},\mathcal{H}}(e\cdot(\epsilon+\delta))\leq m_{\operatorname{PAC},\mathcal{H}}(\epsilon,\delta)\leq O\big{(}m_{\operatorname{Trans},\mathcal{H}}(\epsilon/2)\cdot\log(1/\delta)\big{)}.$
Proof.
Let $m_{\operatorname{Exp},\mathcal{H}}$ denote the sample complexity of learning $\mathcal{H}$ in the expected error regime (i.e., $m_{\operatorname{Exp},\mathcal{H}}(\epsilon)$ equals the number of datapoints needed to incur expected error at most $\epsilon$). Then we have
$m_{\operatorname{Exp},\mathcal{H}}(\epsilon+\delta)\leq m_{\operatorname{PAC},\mathcal{H}}(\epsilon,\delta)\leq O\big{(}m_{\operatorname{Exp},\mathcal{H}}(\epsilon/2)\cdot\log(1/\delta)\big{)}.$
The first inequality follows immediately from the fact that the loss function is bounded above by 1. The second inequality follows from a repetition argument, i.e., a learner attaining expected error $\leq\epsilon/2$ on samples of size $n$ can be boosted to attain expected error $\leq\epsilon$ with probability $\geq 1-\delta$ by using an additional factor of $O(\log(1/\delta))$ many samples, as described in (Daniely and Shalev-Shwartz, 2014).
The first inequality follows inequality from a standard leave-one-out argument of Haussler et al. (1994). The second inequality follows from the fact that for any $S=((x_{1},y_{1}),\ldots,(x_{n},y_{n}))$, there exists an $m_{n}\in\mathbb{N}$ such that $m_{n}$ many independent draws from the uniform distribution over the entries of $S$ has probability at least $\frac{1}{\epsilon}$ of containing exactly $n-1$ elements of $S$, as detailed in (Asilis et al., 2024, Lemma A.1). The claim follows from both of the established chains of inequalities. ∎
We now follow through on porting our results from the transductive model to the PAC model. The following is an immediate consequence of applying Lemma 3.19 to Theorems 3.9 and 3.7.
Corollary 3.20.
Let $\mathcal{X}$ be a domain, $\mathcal{Y}$ a label set, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Suppose that the loss function $\ell$ is bounded to $[0,1]$ and satisfies either of the following conditions:
  • $\ell$ is a metric on $\mathcal{Y}$, or
  • $(\mathcal{Y},d)$ is a compact metric space and $\ell$ is continuous with respect to this topology.
Then if all finite projections of $\mathcal{H}$ are learnable with realizable PAC sample function $m:(0,1)^{2}\to\mathbb{N}$, $\mathcal{H}$ is learnable with sample complexity $O\big{(}m(\frac{\epsilon}{4e},\frac{\epsilon}{4e})\log(1/\delta)\big{)}$.
Let us mention briefly that the connection between transductive learning and PAC learning may not be as tight in the agnostic case as in the realizable case. Through a straightforward use of Markov’s inequality and a repetition argument, one can show that agnostic PAC sample complexities exceed transductive by at most a factor of $1/\epsilon$, but this is an unimpressive bound.